OpenStudy (anonymous):

Write 2 ln 6 - ln 9 as a single natural logarithmic

OpenStudy (anonymous):

ln(36/9)=ln4

OpenStudy (unklerhaukus):

because\[a\ln x=\ln x^a\]\[\ln y-\ln z =\ln\frac yz\]